OpenStudy (xo.mimii):

A company sells boxes of small blue pom-poms (b) for $42 and boxes of large orange pom-poms (p) for $52. They make batches of small blue pom-poms that fill 8 boxes and batches of large orange pom-poms that fill 4 boxes. The company only has 40 boxes. Which system of equations helps the company plan to make $350?

OpenStudy (zachary1):

F(x) = 4x - 5...when x = -3 f(-3) = 4(-3) - 5 f(-3) = -12 - 5 f(-3) = - 17

likeabossssssss (likeabossssssss):

welcome to os